The mean weight (in kg) for 8 adults males are given as follows:

70 72 65 80 75 76 68 78

Construct a 90% confidence interval estimate for the mean weight for all adult males.


1
Expert's answer
2021-11-22T13:24:09-0500
"\bar{x}=\dfrac{70+ 72 +65+ 80+ 75 +76+ 68+78}{8}=73"

"s^2=\dfrac{1}{8-1}((70-73)^2+(72-73)^2+(65-73)^2"

"+(80-73)^2+(75-73)^2+(76-73)^2"

"+(68-73)^2+(78-73)^2)=\dfrac{186}{7}"

"s=\sqrt{s^2}=\sqrt{\dfrac{186}{7}}\approx5.154748"

The critical value for "\alpha = 0.1" and "df = n-1 = 7" degrees of freedom is"t_c = z_{1-\alpha/2; n-1} = 1.894577."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(73-1.894577\times\dfrac{5.154748}{\sqrt{8}},"

"73+1.894577\times\dfrac{5.154748}{\sqrt{8}})"

"=(69.547, 76.453)"

Therefore, based on the data provided, the 90% confidence interval for the population mean is "69.547 < \mu < 76.453," which indicates that we are 90% confident that the true population mean "\mu" is contained by the interval "(69.547, 76.453)."